OpenStudy (anonymous):

Find the remainder when f(x) is divided by (x - k) f(x) = 4x3 - 6x2 + 3x + 1; k= -2 -3 -61 26 -49

OpenStudy (anonymous):

(x-k) = (x- (-2) = (x+2) use reminder theorem

OpenStudy (3psilon):

You can plug in just 2 as x according to the remainder theorem

OpenStudy (anonymous):

f(-2)= 4x3 - 6x2 + 3x + 1

Parth (parthkohli):

Actually, f(-2) = 4(-2)^3 - 6(-2)^2 + 3(-2) + 1

OpenStudy (anonymous):

-2 | 4 -6 3 1 -8 28 -62 4 -14 31 -61 i.e. \[\frac{4x^3-6x^2+3x+1}{x+2}=4x^2-14x+31-\frac{61}{x+2}\]

OpenStudy (anonymous):

so the answer would be -61??